# ThyroidEffi 1.0: A Cost-Effective System for High-Performance Multi-Class Thyroid Carcinoma Classification

### v1

@@ Abstract
Background: Automated classification of thyroid fine needle aspiration biopsy (FNAB) images faces challenges in limited data, inter-observer variability, and computational cost. Efficient, interpretable models are crucial for clinical support. Objective: To develop and externally validate a deep learning system for the multi-class classification of thyroid FNAB images into three key categories that directly guide post-biopsy treatment decisions in Vietnam: benign (B2), suspicious for malignancy (B5), and malignant (B6), while achieving high diagnostic accuracy with low computational overhead. Methods: Our framework features: (1) YOLOv10-based cell cluster detection for informative sub-region extraction and noise reduction; (2) a curriculum learning-inspired protocol sequencing localized crops to full images for multi-scale feature capture; (3) adaptive lightweight EfficientNetB0 (4 millions parameters) selection balancing performance and efficiency; and (4) a Transformer-inspired module for multi-scale, multi-region analysis. External validation used 1,015 independent FNAB images. Results: ThyroidEffi Basic achieved a macro F1 of 89.19% and AUCs of 0.98 (B2), 0.95 (B5), and 0.96 (B6) on the internal test set. External validation yielded AUCs of 0.9495 (B2), 0.7436 (B5), and 0.8396 (B6). ThyroidEffi Premium improved macro F1 to 89.77%. Grad-CAM highlighted key diagnostic regions, confirming interpretability. The system processed 1000 cases in 30 seconds, demonstrating feasibility on widely accessible hardware like a 12-core CPU. Conclusions: This work demonstrates that high-accuracy, interpretable thyroid FNAB image classification is achievable with minimal computational demands.
Keywords Papillary Thyroid Carcinoma  $\cdot$ Medical Image Classification  $\cdot$ Fine-Needle Aspiration Biopsy  $\cdot$ The Bethesda System  $\cdot$ Deep Learning  $\cdot$ Multi-class  $\cdot$ Diagnosis
@@ 1.1 Background and Significance
Thyroid carcinoma is a prevalent endocrine malignancy worldwide [13]. Accurate and efficient diagnosis of thyroid conditions is crucial for effective treatment and improved patient outcomes. Fine needle aspiration biopsy (FNAB) is a minimally invasive procedure widely used for assessing thyroid abnormalities and guiding treatment decisions [1]. Historically, FNAB diagnosis has relied heavily on cytopathologists’ expertise. However, this reliance leads to potential inter-observer variability [3]. The global shortage of trained cytopathologists further emphasizes the need for automated diagnostic methods. Deep learning has shown promise in various medical imaging tasks [25]. Moreover, cost-efficiency is essential for real-world deployment, particularly in resource-constrained settings.
@@ 1.2 Literature Review
Several studies have explored automated classification of thyroid cytology images using deep learning, employing various approaches. Early works often focused on binary classification using Convolutional Neural Networks (CNNs), such as Sanyal et al., [30], which classified papillary thyroid carcinoma (PTC) and non-PTC. These studies demonstrated the potential of deep learning but oversimplified the diagnostic task by ignoring the nuanced categories of TBSRTC. Other studies adopted a Whole Slide Image (WSI) approach, utilizing two-stage CNNs for region of interest identification and subsequent classification, as seen in Dov et al., [7], Elliott Range et al., [11], Duan et al., [9], Assaad et al., [2], and Dov et al., [6]. While WSI analysis allows for comprehensive slide evaluation, it often suffers from increased computational complexity and requires specialized equipment. Another line of research focused on fragment-level analysis, as exemplified by Guan et al., [15] and Duc et al., [10]. These studies achieved high accuracy on fragmented images using models like VGG-16, Inception-v3, and DenseNet161, sometimes combined with techniques like AdaBoost. However, they faced challenges in aggregating fragment-level predictions to obtain reliable patient-level diagnoses and often required manual pre-processing. Some studies explored alternative techniques, such as Correlation Optical Diffraction Tomography (CODT) combined with machine learning algorithms (Lee et al., [23]) or fluorescence polarization imaging (Jermain et al., [22]) for cell segmentation and classification. While innovative, these approaches often require specialized equipment and may not be readily applicable in all clinical settings. More recently, studies have leveraged larger datasets and more sophisticated architectures, such as the ThyroPower system developed by Wang et al., 2024b [37], demonstrating promising results in distinguishing benign from TBSRTC III+ and V+ categories. Ma et al., [28] focused on intraoperative thyroid nodule diagnosis (IOPD) using surgical samples, combining computer vision techniques with CNNs and SVMs.
@@ 1.3 Limitations of Existing Techniques
Despite the extensive research, achieving robust multi-class classification of thyroid cytology images, aligned with the Bethesda System for Reporting Thyroid Cytopathology (TBSRTC), remains a significant challenge [24, 21, 26, 27, 12, 16, 39, 33, 29, 4, 17, 40].
Firstly, many are constrained by small datasets, hindering model generalizability. Secondly, patient-level prediction aggregation remains a non-trivial issue for fragment-based approaches. Thirdly, computational complexity often restricts the real-time applicability of intricate models, particularly those employing whole slide images (WSI) or multi-stage processing. Crucially, a key limitation lies in attaining consistently high multi-class classification performance. While recent investigations (from two research groups, one in China Wang et al., 2024b [37] and one in the United States [8]) have explored multi-class classification, achieving consistently high accuracy across all categories, notably for categories V (Suspicious for Malignancy) and VI (Malignant), continues to be a major hurdle. Details are provided in Figure 1.
@@ 1.4 Research Objectives
This research aims to develop a robust and cost-efficient deep learning model for the multi-class classification of thyroid FNAB slide images (the scope of this research is detailed in Figure 2). Our specific objectives are to: achieve high classification accuracy; maintain low deployment cost (few parameters, low hardware requirements, fast inference time), which is particularly useful for practical implementations in resource-constrained settings like developing countries; and provide good interpretability of diagnostic decisions, enhancing clinical applicability. We focus on Bethesda categories II (benign), V (suspicious for malignancy), and VI (malignant) because these are the most clinically consequential categories for downstream patient management in Vietnam. These three categories directly dictate critical treatment pathways: routine monitoring for Bethesda II, further investigation for Bethesda V, and immediate surgical intervention for Bethesda VI. In contrast, categories I, III, and IV often necessitate repeat FNAB procedures or further molecular testing before a definitive course of action can be determined. Therefore, accurate classification of categories II, V, and VI has the most immediate and substantial impact on patient care, especially as the accurate identification of categories V and VI directly influences the decision to perform surgery or pursue other aggressive treatment options.
@@ 2.1 Dataset
The dataset consists of 1,804 high-resolution microscopic images obtained from fine-needle aspiration biopsy (FNAB) samples, each corresponding to a unique patient. These images were acquired at the 108 Military Central Hospital between 2018 and July 2024 using an Olympus BX43 optical microscope equipped with a high-resolution digital camera and a specific Diff-Quick staining protocol. The exclusive use of a 40x magnification and a standardized Diff-Quick staining protocol aimed to minimize inter-institutional variability in image acquisition. The dataset includes 482 benign (B2), 541 suspected malignant (B5), and 871 malignant (B6) samples, reflecting the real-world distribution often encountered in clinical settings, with Bethesda VI (Malignant) cases being more prevalent. The selection of these three classes (Bethesda II, V, and VI) was based on their clinical significance and diagnostic challenge, as categories V and VI often require careful differentiation to guide appropriate patient management.
@@ 2.2 Methodological
To address these challenges while achieving our stated objectives, we designed a novel four-module framework (M1 to M4), detailed visually in Figure 3. These four seamlessly integrated modules work synergistically to mitigate the aforementioned limitations in the following ways:
@@ 2.2.2 New Training Protocol (M2)
This training protocol builds upon the principle that patterns learned last are better retained and more effective in neural networks or human cognition. Thus, placing patient-level samples (Set A) at the end of each epoch enhances learning efficacy. By strategically prioritizing the learning of full-image context towards the end of each epoch, the model is encouraged to consolidate its understanding of the overall diagnostic landscape, ultimately improving its ability to make accurate patient-level classifications. A notable feature of our training approach is the structured sequence of images within each epoch, adhering to a defined data augmentation hierarchy. Images are processed sequentially from sets E, D, C, B, to A, as shown in figure 3, M2. This arrangement enables the model to progressively learn features at multiple scales, beginning with high-noise images (set E) and advancing to low-noise sets. The specific order of images in the training sequence (E to A) is designed to gradually increase the complexity and contextual richness of the training data. By starting with smaller, high-noise regions, the model is encouraged to focus on identifying fundamental cellular features. As the training progresses through sets D, C, and B, the model is exposed to increasingly larger and more complete image regions, allowing it to integrate these local features into a broader understanding of the overall cytological context.
@@ 2.2.3 Model Selection Strategy (M3)
The selected models span various complexities, from highly parameterized deep networks to lightweight versions of the same architecture. This range allows for an in-depth evaluation of trade-offs in terms of computational efficiency, model complexity, and performance accuracy, particularly in handling the challenges of imbalanced datasets in multi-class classification. The models considered include traditional convolutional neural networks (CNNs) such as VGG [32], ResNet [18], MobileNet [19], DenseNet [20], EfficientNet [34], as well as Transformer-based architectures like Vision Transformer (ViT) [5], offering a broad spectrum for comparison. Additionally, customized variants with adjusted loss functions to address class imbalance Appendix E. Validation, such as class weighting techniques, were explored. This adjustment is crucial to mitigate model bias toward majority classes, a common challenge in imbalanced datasets.
This evaluation framework is designed to provide critical insights into which model architectures and configurations deliver the best balance between computational efficiency and classification accuracy, particularly when dealing with complex, imbalanced data. The findings will help guide the selection of a model that not only meets performance benchmarks but also aligns with resource constraints common in clinical and real-time diagnostic settings.
@@ 3 Results
In response to the reviewers’ request for a comparison with previous research methodologies applied to our dataset, we note that studies in this field generally lack direct comparisons with one or more prior implementations. This stems from several key differences:
@@ 3.1 ThyroidEffi Basic Model Performance on Test Set
Table 2 shows a comparison of macro F1 scores across 24 experiments with different models and data augmentation configurations. The ×$\times 34$ data augmentation approach led to substantial performance improvements across most models, except for ViT-B16 (highlighted in red). EfficientNetB0, trained with the augmented dataset, achieved the highest macro F1 score of 89.19% and is designated as ThyroidEffi Basic.
@@ 3.2 Comparison of ThyroidEffi Basic and ThyroidEffi Premium
True Label B2 B5 B6 Total
B2 57 2 2 61
B5 0 85 11 96
B6 2 15 98 115
@@
The ThyroidEffi Basic model also achieved AUC scores of B2: 0.98, B5: 0.95, and B6: 0.96 on the test set.
@@ 3.3 Real-World Application Results
  • Area under the curve (AUC): B2: 0.9495, B5: 0.7436, B6: 0.8396.
  • The classification report and confusion matrix for the real-world application data are shown in table 4 and table 5, respectively.
@@
B2 B5 B6 Macro Avg
@@
True Label B2 B5 B6 Total
B2 261 30 9 300
B5 44 143 128 315
B6 16 92 292 400
@@ 4.3 Model Explainability and Visualization
Figure 4: Visualization of data distribution before and after applying the ThyroidEffi Basic model for classifying B2, B5, and B6 categories:
- Top row: Original dataset reduced to 2D using PCA [14], t-SNE, and UMAP [38], showing overlapping clusters and poor separability among classes.
- Bottom row: Dataset transformed by ThyroidEffi Basic into a 3-dimensional latent space (representing probabilities for B2, B5, B6) and reduced to 2D. Clear cluster separation highlights the model’s ability to extract discriminative features.
Figure 4 visually demonstrates the effectiveness of the ThyroidEffi Basic model in transforming complex image features into a more discriminative latent space. The clear separation of clusters in the transformed space highlights the model’s ability to extract relevant features for classifying B2, B5, and B6 categories. This visualization provides insights into the model’s decision-making process and supports its validity.
Figure 5: CAM of a slide image (left) and 12 sub-images (right) after pass ThyroidEffi Basic

### v2

@@ Abstract
Background: Automated classification of thyroid Fine Needle Aspiration Biopsy (FNAB) images faces challenges in limited data, inter-observer variability, and computational cost. Efficient, interpretable models are crucial for clinical support. Objective: To develop and externally validate a deep learning system for the multi-class classification of thyroid FNAB images into three key categories that directly guide post-biopsy treatment decisions in Vietnam: Bethesda II: Observation / Non-Surgical Management (Benign), Bethesda I, III, IV, V: Further Investigation / Consideration for Intervention (Indeterminate/Suspicious), and Bethesda VI: Surgical Intervention (Malignant), while achieving high diagnostic accuracy with low computational overhead. Methods: Our pipeline features: (1) YOLOv10-based cell cluster detection for informative sub-region extraction and noise reduction; (2) a curriculum learning-inspired protocol sequencing localized crops to full images for multi-scale feature capture; (3) adaptive lightweight EfficientNetB0 (4 millions parameters) selection balancing performance and efficiency; and (4) a Transformer-inspired module for multi-scale, multi-region analysis. External validation used 1,015 independent FNAB images. Results: ThyroidEffi Basic achieved a macro F1 of 89.19% and Area Under the Curve (AUC) s of 0.98 (Benign), 0.95 (Indeterminate/Suspicious), and 0.96 (Malignant) on the internal test set. External validation yielded AUC s of 0.9495 (Benign), 0.7436 (Indeterminate/Suspicious), and 0.8396 (Malignant). ThyroidEffi Premium improved macro F1 to 89.77%. Gradient-weighted Class Activation Mapping (Grad-CAM) highlighted key diagnostic regions, confirming interpretability. The system processed 1000 cases in 30 seconds, demonstrating feasibility on widely accessible hardware like a 12-core CPU. Conclusions: This work demonstrates that high-accuracy, interpretable thyroid FNAB image classification is achievable with minimal computational demands.
Keywords Thyroid Carcinoma  $\cdot$ Fine-Needle Aspiration Biopsy  $\cdot$ Medical Image Classification  $\cdot$ Deep Learning  $\cdot$ Multi-class Classification  $\cdot$ Bethesda System  $\cdot$ Cost-Effective
Graphical Abstract
Figure 1: Graphical Abstract
@@ 1.1 Background and Significance
Thyroid carcinoma is a prevalent endocrine malignancy worldwide [1]. Accurate and efficient diagnosis of thyroid conditions is crucial for effective treatment and improved patient outcomes. FNAB is a minimally invasive procedure widely used for assessing thyroid abnormalities and guiding treatment decisions [2]. Historically, FNAB diagnosis has relied heavily on cytopathologists’ expertise. However, this reliance leads to potential inter-observer variability [3]. The global shortage of trained cytopathologists further emphasizes the need for automated diagnostic methods. Deep learning has shown promise in various medical imaging tasks [4]. Moreover, cost-efficiency is essential for real-world deployment, particularly in resource-constrained settings.
@@ 1.2 Literature Review
Several studies have explored automated classification of thyroid cytology images using deep learning, employing various approaches. Early works often focused on binary classification using Convolutional Neural Network (CNN), such as Sanyal et al. [5], which classified Papillary Thyroid Carcinoma (PTC) and non-PTC. These studies demonstrated the potential of deep learning but oversimplified the diagnostic task by ignoring the nuanced categories of The Bethesda System for Reporting Thyroid Cytopathology (TBSRTC). Other studies adopted a Whole Slide Image (WSI) approach, utilizing two-stage CNN s for Region of Interest (ROI) identification and subsequent classification, as seen in Dov et al. [6], Elliott Range et al. [7], Duan et al. [8], Assaad et al. [9], and Dov et al. [10]. While WSI analysis allows for comprehensive slide evaluation, it often suffers from increased computational complexity and requires specialized equipment. Another line of research focused on fragment-level analysis, as exemplified by Guan et al. [11] and Duc et al. [12]. These studies achieved high accuracy on fragmented images using models like VGG-16, Inception-v3, and DenseNet161, sometimes combined with techniques like AdaBoost. However, they faced challenges in aggregating fragment-level predictions to obtain reliable patient-level diagnoses and often required manual pre-processing. Some studies explored alternative techniques, such as Correlation Optical Diffraction Tomography (CODT) combined with machine learning algorithms (Lee et al. [13]) or fluorescence polarization imaging (Jermain et al. [14]) for cell segmentation and classification. While innovative, these approaches often require specialized equipment and may not be readily applicable in all clinical settings. More recently, studies have leveraged larger datasets and more sophisticated architectures, such as the ThyroPower system developed by Wang et al. [15], demonstrating promising results in distinguishing benign from TBSRTC III+ and V+ categories. Ma et al. [16] focused on Intraoperative Thyroid Nodule Diagnosis (IOPD) using surgical samples, combining computer vision techniques with CNN and Support Vector Machine (SVM).
@@ 1.3 Limitations of Existing Techniques
Despite the extensive research, achieving robust multi-class classification of thyroid cytology images, aligned with TBSRTC, remains a significant challenge [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
Firstly, many are constrained by small datasets, hindering model generalizability. Secondly, patient-level prediction aggregation remains a non-trivial issue for fragment-based approaches. Thirdly, computational complexity often restricts the real-time applicability of intricate models, particularly those employing WSI or multi-stage processing. Crucially, a key limitation lies in attaining consistently high multi-class classification performance. While recent investigations (from two research groups, one in China Wang et al. [15] and one in the United States [6, 29]) have explored multi-class classification, achieving consistently high accuracy across all categories, notably for categories V (Suspicious for Malignancy) and VI (Malignant), continues to be a major hurdle. Details are provided in Figure 2.
@@ 1.4 Research Objectives
This research aims to develop a robust and cost-efficient deep learning model for the classification of thyroid nodule FNAB slide images (the scope of this research is detailed in Figure 3). Our specific objectives are to achieve high classification accuracy; maintain low deployment cost (few parameters, low hardware requirements, fast inference time), which is particularly useful for practical implementations in resource-constrained settings like developing countries; and provide good interpretability of diagnostic decisions, enhancing clinical applicability.
TBSRTC classifies thyroid nodule FNAB results into six distinct categories (Bethesda I-VI). However, in clinical practice (especially in Vietnam), the interpretation of these six cytological categories ultimately guides patient management into one of three primary treatment pathways: Observation/Non-Surgical Management (typically corresponding to Bethesda II), Further Investigation/Consideration for Intervention (encompassing Bethesda I, III, IV, and V), or Surgical Intervention (primarily corresponding to Bethesda VI).
Our model is specifically designed to classify FNAB images directly into these three clinically relevant groups: Benign (corresponding to Bethesda II), Indeterminate/Suspicious (corresponding to the combined Bethesda I, III, IV, and V categories), and Malignant (corresponding to Bethesda VI). This approach is motivated by the direct link between these three groups and the most crucial decisions in patient care. Accurate classification into these groups provides crucial support for clinicians in determining the appropriate primary treatment pathway for a patient, streamlining the diagnostic process.
Compared to previous studies, which often focused solely on binary classification between benign and malignant, thus overlooking the crucial indeterminate and suspicious categories, or which addressed multi-class classification based on TBSRTC but did not cover all six categories or evaluate performance comprehensively across all included classes, our research offers a more clinically aligned classification strategy. By training our model to classify directly into the three main management groups that encompass the entire Bethesda spectrum, this study provides a more pragmatic solution directly supporting the established clinical workflow, particularly beneficial for guiding treatment decisions based on current image data in resource-constrained settings.
@@ 2.1 Dataset
The dataset consists of 1,804 high-resolution microscopic images obtained from FNAB samples, each corresponding to a unique patient. These images were acquired at the 108 Military Central Hospital between 2018 and July 2024 using an Olympus BX43 optical microscope equipped with a high-resolution digital camera and a specific Diff-Quick staining protocol. The exclusive use of a 40x magnification and a standardized Diff-Quick staining protocol aimed to minimize inter-institutional variability in image acquisition. The dataset includes 482 (Benign), 541 (Indeterminate/Suspicious), and 871 (Malignant) samples, reflecting the real-world distribution often encountered in clinical settings, with Bethesda VI (Malignant) cases being more prevalent. The selection of these three classes (Benign, Indeterminate/Suspicious, Malignant) was based on their clinical significance and diagnostic challenge, as accurately distinguishing between these groups is paramount for directing patients towards appropriate management pathways, ranging from observation to immediate surgical intervention.
@@ 2.2 Methodological
To address these challenges while achieving our stated objectives, we designed a novel four-module pipeline (M1 to M4), detailed visually in Figure 4. These four seamlessly integrated modules work synergistically to mitigate the aforementioned limitations in the following ways:
@@ 2.2.2 New Training Protocol (M2)
This training protocol builds upon the principle that patterns learned last are better retained and more effective in neural networks or human cognition. Thus, placing patient-level samples (Set A) at the end of each epoch enhances learning efficacy. By strategically prioritizing the learning of full-image context towards the end of each epoch, the model is encouraged to consolidate its understanding of the overall diagnostic landscape, ultimately improving its ability to make accurate patient-level classifications.
During the training process, each individual augmented image or cropped region generated through the M1 strategy (from sets A, B, C, D, and E) was treated as a separate training sample. These samples were assigned the overall diagnostic label (Benign, Indeterminate/Suspicious, or Malignant) of their original full image. A notable feature of our training approach is the structured sequence of images within each epoch, adhering to a defined data augmentation hierarchy. Images are processed sequentially from sets E, D, C, B, to A, as shown in figure 4, M2. This arrangement enables the model to progressively learn features at multiple scales, beginning with high-noise images (set E) and advancing to low-noise sets. The specific order of images in the training sequence (E to A) is designed to gradually increase the complexity and contextual richness of the training data. By starting with smaller, high-noise regions, the model is encouraged to focus on identifying fundamental cellular features. As the training progresses through sets D, C, and B, the model is exposed to increasingly larger and more complete image regions, allowing it to integrate these local features into a broader understanding of the overall cytological context.
@@
While individual cropped regions from sets C, D, and E may not encompass all diagnostic features of the full image, assigning the full image’s label to these localized views is a deliberate strategy. This approach, combined with the curriculum learning sequence of M2 (progressing from localized crops to full images), trains the model to identify crucial cellular and tissue patterns even in limited views, while the subsequent training on broader contexts (sets B and A) allows the model to integrate these local findings for accurate patient-level classification, effectively mitigating the risk of misinterpretation based solely on partial information. Especially when combined with M4, this allows for the integration of both small local image features (which are often lost if images are resized to smaller dimensions to reduce deep learning model size) and global, large-scale features (which are often lost if only selected regions are cropped from the image). Ultimately, learning both local and global level features in M2 and combining them in M4 leads to better prediction quality by utilizing more feature information.
@@ 2.2.3 Model Selection Strategy (M3)
The selected models span various complexities, from highly parameterized deep networks to lightweight versions of the same architecture. This range allows for an in-depth evaluation of trade-offs in terms of computational efficiency, model complexity, and performance accuracy, particularly in handling the challenges of imbalanced datasets in multi-class classification. The models considered include traditional CNN such as VGG [31], ResNet [32], MobileNet [33], DenseNet [34], EfficientNet [35], as well as Transformer-based architectures like Vision Transformer (ViT) [36], offering a broad spectrum for comparison. Additionally, customized variants with adjusted loss functions to address class imbalance Appendix E. Validation, such as class weighting techniques, were explored. This adjustment is crucial to mitigate model bias toward majority classes, a common challenge in imbalanced datasets.
This evaluation is designed to provide critical insights into which model architectures and configurations deliver the best balance between computational efficiency and classification accuracy, particularly when dealing with complex, imbalanced data. The findings will help guide the selection of a model that not only meets performance benchmarks but also aligns with resource constraints common in clinical and real-time diagnostic settings.
@@ 3 Results
When comparing with previous research methods applied to our dataset, we note that studies in this area often lack direct comparison with one or more previous implementations. This stems from several key differences:
@@ 3.1 ThyroidEffi Basic Model Performance on Test Set
Table 2 shows a comparison of macro F1 scores. We prioritized using the macro F1 score because it provides a balanced measure, ensuring accurate performance assessment across all three classification classes, and is less sensitive to data imbalance compared to overall accuracy, thus providing a more reliable representation of the model’s overall effectiveness for all classes. The table compares the performance of common CNN models replicated from previous studies (shown in column 2) with the performance of these models when implemented within our pipeline (shown in column 3). This comparison, excluding ViT (which is a new model architecture included due to its prevalence and performance in recent studies on large datasets), highlights the impact of our pipeline and training protocol on model performance. The ×$\times 34$ data augmentation approach led to substantial performance improvements across most models, except for ViT-B16 (highlighted in red). EfficientNetB0, trained with the augmented dataset, achieved the highest macro F1 score of 89.19% and is designated as ThyroidEffi Basic.
@@ 3.2 Comparison of ThyroidEffi Basic and ThyroidEffi Premium
True Label Benign Indeterminate/Suspicious Malignant Total
Benign 57 2 2 61
Indeterminate/Suspicious 0 85 11 96
Malignant 2 15 98 115
@@
The ThyroidEffi Basic model also achieved AUC scores of Benign: 0.98, Indeterminate/Suspicious: 0.95, and Malignant: 0.96 on the test set.
@@ 3.3 Real-World Application Results
Benign Indeterminate/Suspicious Malignant Macro Avg
@@
True Label Benign Indeterminate/Suspicious Malignant Total
Benign 261 30 9 300
Indeterminate/Suspicious 44 143 128 315
Malignant 16 92 292 400
@@ 4.1 Cost-Effectiveness and High Performance of the System
It is important to clarify that during inference and real-world deployment, the model receives only the raw or minimally pre-processed image data as input. It does not have access to or receive any pre-existing information about the pathologist’s ground truth diagnosis or the original Bethesda category (I-VI) of the image. The model’s output is solely its prediction of the likelihood distribution across the three defined clinical management groups based on the learned features from the image.
@@ 4.3 Model Explainability and Visualization
Figure 5: Visualization of data distribution before and after applying the ThyroidEffi Basic model for classifying Benign, Indeterminate/Suspicious, and Malignant categories:
- Top row: Original dataset reduced to 2D using Principal Component Analysis (PCA) [38], t-distributed Stochastic Neighbor Embedding (t-SNE), and Uniform Manifold Approximation and Projection (UMAP) [39], showing overlapping clusters and poor separability among classes.
- Bottom row: Dataset transformed by ThyroidEffi Basic into a 3-dimensional latent space (representing probabilities for Benign, Indeterminate/Suspicious, Malignant) and reduced to 2D. Clear cluster separation highlights the model’s ability to extract discriminative features.
Figure 5 visually demonstrates the effectiveness of the ThyroidEffi Basic model in transforming complex image features into a more discriminative latent space. The clear separation of clusters in the transformed space highlights the model’s ability to extract relevant features for classifying Benign, Indeterminate/Suspicious, and Malignant categories. This visualization provides insights into the model’s decision-making process and supports its validity.
Figure 6: Grad-CAM of a slide image (left) and 12 sub-images (right) after pass ThyroidEffi Basic